OpenStudy (anonymous):

A small publishing company is planning to publish a new book. The production costs will include one-time fixed costs (such as editing) and variable costs (such as printing). The one-time fixed costs will total $21,021. The variable costs will be $12.25 per book. The publisher will sell the finished product to bookstores at a price of $20.50 per book. How many books must the publisher produce and sell so that the production costs will equal the money from sales?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

if you publish say \(n\) books then the total cost will be \[12.25n+21021\] right?

OpenStudy (misty1212):

this is where you say "yes of course" or "how did you get that?"

OpenStudy (anonymous):

yes but how did you get that ?

OpenStudy (misty1212):

the fixed cost is fixed, you have to pay it no matter how many you print

OpenStudy (misty1212):

if you print one book you pay the fixed cost plus 12.25 so \[12.25+21021\] if you print two books you have to pay two times 12.25 and the fixed cost so \[12.25\times 2+21021\]

OpenStudy (misty1212):

and if you print \(n\) books you have to pay \[12.25\times n+21021\]

OpenStudy (misty1212):

now for each book you sell, you make 20.50 so for n books you make \[20.50n\] dollars

OpenStudy (misty1212):

you want to know where they are equal so set \[12.25n+21021=20.50n\] and solve for \(n\)